# Cold Jupiters and small planets: friends, foes, or indifferent?

### v1

@@ 4.1 Frequency of CJs at average stellar metallicity and mass
At the average stellar metallicity ($\overline{\rm[Fe/H]}=-0.011\pm 0.005$) and mass ($\overline{M_{\star}}=0.916\pm 0.012\,\hbox{$\mathrm{M}_{\odot}$}$) of the SP (transit+RV) stars we determined $f_{\rm CJ|SP}=11.1^{+2.5}_{-1.8}\%$ for the CJ mass range $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$. Adding the 25 mixed systems yielded $f_{\rm CJ|SP+Mixed}=12.5^{+2.5}_{-1.8}\%$. Going back to the original sample of BL24 with $\overline{\rm[Fe/H]}=-0.001\pm 0.005$ and $\overline{M_{\star}}=0.923\pm 0.014\,\hbox{$\mathrm{M}_{\odot}$}$, which includes the “mixed systems”, and accounting for its average completeness [^8] This was extracted from Fig. A.1 in BL24., we found a fully compatible result: $15.2^{+3.1}_{-2.3}\%$. It is interesting to note that both occurrence rates well agree within $1\sigma$ with the estimate of $f_{\rm CJ|SP}=8.7^{+7.3}_{-2.7}\%$ by B23, whose smaller sample is on average slightly poorer in metals by $\sim 0.06$ dex, and are lower than previous claims of $f_{\rm CJ|SP}\sim 40\%$ (Sect. 1).
As for the comparison sample, at its average metallicity ($\overline{\rm[Fe/H]}=-0.072\pm 0.009$) and mass ($\overline{M_{\star}}=0.994\pm 0.004\,\hbox{$\mathrm{M}_{\odot}$}$), we found $f_{\rm CJ}=9.8^{+0.9}_{-0.8}\%$, which is fully compatible with $f_{\rm CJ|SP}$ [^9] The possible, though not significant, hint for a SP-CJ anti-correlation reported by B23 can be ruled out as our $f_{\rm CJ}$ recomputed from the AAT survey is approximately half that derived from Wittenmyer et al. (2020)..
From our occurrence rates $f_{\rm CJ|SP}$ and $f_{\rm CJ}$, and the frequency of SPs $f_{\rm SP}=28.1^{+6.6}_{-5.1}\%$ from Rosenthal et al. (2022) (see also Zhu et al. 2018), we can derive the conditional probability that a solar-type star with a CJ also hosts a SP (Zhu & Wu 2018), namely $f_{\rm SP|CJ}=f_{\rm SP}\times f_{\rm CJ|SP}\leavevmode\nobreak\ /\leavevmode\nobreak\ f_{\rm CJ}=32\pm 11\%$. This agrees within $1\sigma$ with the $f_{\rm SP|CJ}$ computed from a sample of 115 solar-type stars known to host a CJ (Ruggieri, A., et al. in prep.), and is substantially lower than the estimate of $f_{\rm SP|CJ}\sim 100\%$ in Bryan et al. (2019).
@@ 4.2 Frequency of CJs as a function of stellar metallicity
We report $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ for the three sub-intervals of stellar metallicity (Sect. 3) in Table 19, and display them in Fig. 1 (see also Fig. 9 for $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$). For the same metallicity bins, we also give the estimates of $f_{\rm CJ}$ from the AAT, CLS, and HARPS surveys separately in Table 23, which are fully consistent with each other (see Fig. 8).
By comparing $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ in Table 19, we found that (i) at sub-solar and solar metallicities, they are fully compatible within $1\sigma$, and (ii) at super-solar metallicity, the value of $f_{\rm CJ|SP}$ is higher than $f_{\rm CJ}$ by approximately $3/2$ ($\sim 30\%$ vs $\sim 20\%$; Fig. 1), but the significance of this difference is low, that is $1.7\sigma$ (compared to $2.7\sigma$ in BL24 for $\rm[Fe/H]>0$).
Figure 1: Occurrence rates of cold Jupiters with $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ in the small planet sample (red solid line) and the merged RV comparison sample (black dashed line) at sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicity. See also Table 19.
Figure 2: Same as Fig. 1 for three different bins of stellar mass, namely $0.6-0.8\,\hbox{$\mathrm{M}_{\odot}$}$ (left), $0.8-1.0\,\hbox{$\mathrm{M}_{\odot}$}$ (center), and $1.0-1.2\,\hbox{$\mathrm{M}_{\odot}$}$ (right); see Table 20.
@@ 4.3 Frequency of CJs as a function of stellar metallicity and mass
We report $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ for the same bins of stellar metallicity as above, and for the three ranges of stellar mass discussed in Sect. 3, in Table 20. We show them in Fig. 2 (see also Fig. 10 for $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$).
We found an excess of CJs in SP systems at $2.5\sigma$ significance level only at super-solar mass and metallicity, namely $M_{\star}\geq 1.0\,\hbox{$\mathrm{M}_{\odot}$}$ and $\rm[Fe/H]>0$ (see the right plot in Fig. 2, and Table 20), in agreement with the recent findings of BL25. This is the origin of the higher $f_{\rm CJ|SP}$ compared to $f_{\rm CJ}$ at super-solar metallicity and integrated stellar masses (Sect- 4.2 and Fig. 1).
@@ 4.4 SP-CJ relations as a function of SP and CJ multiplicity
We explored possible SP-CJ trends as a function of SP multiplicity, for which purpose we chose the $0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ mass range (instead of $0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$), given the higher number of systems with both CJs and SPs (27 compared to 23). Indeed, if the presence of CJs really hindered the formation/migration of SPs, a lower degree of SP multiplicity would be expected in the presence of CJs, for instance because low-mass “jumpers” overtaking the CJ towards inner orbits most likely form a single SP system (see Fig. 3 in Izidoro et al. 2015). By comparing the frequency of multiple systems with a number of SPs $n_{\rm p}\geq 2$ in the SP sub-samples with and without CJs, we found $f_{n_{\rm p}\geq 2,\rm w/\,CJ}=37^{+10}_{-8}\%$ (10 out of 27 systems) and $f_{n_{\rm p}\geq 2,\rm w/o\,CJ}=43^{+4}_{-3}\%$ (81 out of 190 systems), respectively; these rates are statistically consistent within the uncertainties.
We also searched for possible (anti)correlations as a function of CJ multiplicity to investigate whether two (or more) CJs may hinder the formation of SPs more than a single CJ system (Sect. 1). To this end, we compared the frequency of single SPs in systems with one CJ to that in systems with two (or more) CJs; indeed, the latter should be higher if two (or more) CJs had a more deleterious effect than a single CJ on the formation of multiple SP systems. We did not find any evidence for that: the rates of single SP systems with one CJ and two CJs are compatible, namely $63^{+9}_{-12}\%$ (12 out of 19 systems) and $62^{+13}_{-18}\%$ (5 out of 8 systems), respectively (see Fig. 11).
@@ 4.5 SP-CJ relation as a function of SP composition
We investigated the possible presence of an architecture-composition link by determining the frequency of rocky SPs [^10] We define a rocky planet as a planet that is located between the 100% iron and 100% silicate compositions in the mass-radius diagram of small planets (e.g., Zeng & Sasselov 2013). in CJ systems. This is intended to test whether the formation of rocky planets may be favored in the presence of CJs in the event CJs block the migration of icy sub-Neptunes from beyond the water snowline (Schlecker et al. 2021). We derived $f_{\rm rocky,w/\,CJ}=33^{+13}_{-9}\%$ from 5 transiting systems with rocky planets (HD219134, Kepler-139, Kepler-407, K2-312, K2-364) out of 15 in total (including 3 RV systems that were later found transiting). This shows that systems with both rocky planets and CJs do not represent the majority of the SP-CJ transiting systems.
@@ 4.6 SP multiplicity as a function of CJ eccentricity
When ordering the SP systems with CJs as a function of the CJ eccentricity (Fig. 3), only single or, more rarely, two SPs are found with eccentric ($e\gtrsim 0.4$) CJs, the most extreme case being K2-312/HD 80653 (B23). On the contrary, higher multiplicity (up to $n_{\rm p}=4$ for HD 219134) SP systems are seen in the company of CJs in quasi-circular orbits. This is in line with the expectation of stronger dynamical perturbations at inner orbital separations by eccentric CJs (e.g., Huang et al. 2017).
Figure 3: The architectures of the 27 systems with both small planets and cold Jupiters in the mass range $0.3\leq{\rm M_{p}}\leq 13\,\hbox{$\mathrm{M}_{\rm Jup}$}$, ordered according to increasing orbital eccentricity of the cold Jupiters (from top to bottom; in case of multiple cold Jupiters, the highest eccentricity was considered). The planet circles have sizes proportional to their mass (see legend in the top right corner), and colors depending on the orbital eccentricity (more eccentric planets are shown in yellow). The host stars are color coded as a function of stellar metallicity (metallicity increases from blue to red). All the systems with three or more inner small planets are accompanied by cold Jupiters with low eccentricities.
@@ 5 Discussion and conclusions
5.1 Cold Jupiters and small planets are friends, foes, or indifferent?
@@
In the present work we first investigated the possible SP-CJ relation as a function of stellar metallicity and mass, and did not reveal any correlation or anticorrelation (i.e., SPs and CJs are neither friends nor foes) in eight of the nine mass and metallicity bins, that is at sub-solar mass for any metallicity as well as at super-solar mass for solar and sub-solar metallicity. We instead found an enhancement of CJs in SP systems (i.e., SPs and CJs could be friends) at super-solar mass and metallicity, though at less than $3\sigma$, in general agreement with BL25.
@@
We could not unveil any SP-CJ trends with planet multiplicity and composition, which would further disfavor the anticorrelation models predicting a dearth of CJs in SP systems (Sect. 1). Nonetheless, it should be noted that (i) the limited sub-samples of SP systems with CJs (27 systems) inevitably result in large error bars on the CJ occurrence rates, making it difficult to reveal possible weak correlations or anticorrelations with planet properties; and (ii) the SP multiplicity considered in this work relies on the known (that is, detected) SPs and may be different from the true one, especially if non-transiting SPs with relatively long period ($P\sim 25-100$ d) and/or low mass went undetected by the RV follow up.
@@
Our investigation of the architecture-composition relation in Sect. 4.5 must be considered with caution because it only concerned rocky planets. Yet, in principle it may also apply to rocky planets with tenuous H/He atmospheres, if some available gas was accreted by the super-Earths at the time of their formation, and was not later on stripped away by post-formation processes such as atmospheric photoevaporation (e.g., Lopez & Fortney 2014) or core-powered mass loss (e.g., Ginzburg et al. 2018). However, the well-known degeneracy in SP composition from the measurement of the bulk density alone does not allow us to distinguish whether or which non-rocky planets in the 10 transiting systems with CJs of our SP sample are actually dry (rocky worlds shrouded by thin H/He atmospheres, which formed within the water snowline) or icy (water worlds with possible H/He envelops, which formed beyond the water snowline and then migrated inwards). This makes the investigation of the architecture-composition link challenging.
@@
We showed that higher multiplicity SP systems are found in systems with CJs in low-eccentricity orbits, being immune to dynamical instabilities generated by highly eccentric CJs. In this sense, SPs are not indifferent to CJs, even in the great majority of cases where $f_{\rm CJ|SP}\sim f_{\rm CJ}$.
5.2 Caveats of our, BL24, and BL25 analyses
@@
The occurrence rates derived in this work, BL24, and BL25 may be slightly underestimated mainly because of two reasons, which however should affect both the SP and comparison samples approximately in the same way, making the search for correlations in principle reliable. Firstly, the completenesses computed in our and BL24 works are optimistic because they only account for the formal RV uncertainties, and not for the uncorrelated jitter (as properly done by B23), which in some cases may be considerably higher than the former [^11] Nevertheless, proper Markov chain Monte Carlo analyses of the RV datasets of $\sim 1350$ systems considered here, similarly to B23, would require unrealistic computation times.. Secondly, a few linear or quadratic RV trends, which might later result in CJ signals, were not taken into account in the computation of the occurrence rates, which relies on confirmed CJs only. Nevertheless, among the targets in the SP sample with relatively short temporal baselines, that is $1\leq\Delta T\leq 4$ yr, we noticed only two targets with RV long-term trends still compatible with CJs (namely TOI-1279 and TOI-1778). Moreover, such trends are very often induced by brown dwarfs or low-mass stellar companions, or they correlate with activity indicators (see B23).
@@
Conversely, occurrence rates might decrease if the minimum masses ($M_{\rm p}\sin{i}$) of some of the considered CJs turned into true mass ($M_{\rm p}$) measurements above the 20 (or 13) $\mathrm{M}_{\rm Jup}$ upper limit, corresponding to brown dwarfs or low-mass stellar objects, due to low orbital inclinations (e.g, Xiao et al. 2023).

### v2

@@ 4.1 Frequency of CJs at average stellar metallicity and mass
At the average stellar metallicity ($\overline{\rm[Fe/H]}=-0.011\pm 0.005$) and mass ($\overline{M_{\star}}=0.916\pm 0.012\,\hbox{$\mathrm{M}_{\odot}$}$) of the SP (transit+RV) stars, we determine $f_{\rm CJ|SP}=11.1^{+2.5}_{-1.8}\%$ for the CJ mass range $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$. Including the 25 mixed systems yields $f_{\rm CJ|SP+Mixed}=12.5^{+2.5}_{-1.8}\%$. Returning to the original sample of BL24, which has $\overline{\rm[Fe/H]}=-0.001\pm 0.005$ and $\overline{M_{\star}}=0.923\pm 0.014\,\hbox{$\mathrm{M}_{\odot}$}$ and includes the mixed systems, and accounting for its average completeness [^8] This was extracted from Fig. A.1 in BL24., we find a fully consistent result: $15.2^{+3.1}_{-2.3}\%$. Both occurrence rates are lower than previous claims of $f_{\rm CJ|SP}\sim 40\%$ (Sect. 1). They agree within $1\sigma$ with the estimate of $f_{\rm CJ|SP}=8.7^{+7.3}_{-2.7}\%$ by B23, whose smaller sample is on average slightly more metal-poor by $\sim 0.06$ dex.
For the comparison sample, at its average metallicity ($\overline{\rm[Fe/H]}=-0.072\pm 0.009$) and mass ($\overline{M_{\star}}=0.994\pm 0.004\,\hbox{$\mathrm{M}_{\odot}$}$), we find $f_{\rm CJ}=9.8^{+0.9}_{-0.8}\%$, which is fully compatible with $f_{\rm CJ|SP}$ [^9] The possible, though not significant, hint for an SP-CJ anticorrelation reported by B23 can be ruled out as our $f_{\rm CJ}$ recomputed from the AAT survey is approximately half that derived from Wittenmyer et al. (2020)..
From our occurrence rates $f_{\rm CJ|SP}$ and $f_{\rm CJ}$, and the frequency of SPs $f_{\rm SP}=28.1^{+6.6}_{-5.1}\%$ from Rosenthal et al. (2022) (see also Zhu et al. 2018), we can derive the conditional probability that a solar-type star with a CJ also hosts a SP (Zhu & Wu 2018), namely $f_{\rm SP|CJ}=f_{\rm SP}\times f_{\rm CJ|SP}\leavevmode\nobreak\ /\leavevmode\nobreak\ f_{\rm CJ}=32\pm 11\%$. This agrees within $1\sigma$ with the $f_{\rm SP|CJ}$ computed from a sample of 115 solar-type stars known to host a CJ (Ruggieri et al., in prep.), and is substantially lower than the estimate of $f_{\rm SP|CJ}\sim 100\%$ in Bryan et al. (2019).
@@ 4.2 Frequency of CJs as a function of stellar metallicity
We report $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ for the three subintervals of stellar metallicity (Sect. 3) in Table 13and display them in Fig. 1 (see also Fig. 9 for $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$). For the same metallicity bins, we also provide the estimates of $f_{\rm CJ}$ from the AAT, CLS, and HARPS surveys separately in Table 17; these are fully consistent with each other (see Fig. 8).
By comparing $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ in Table 13, we find that (i) at sub-solar and solar metallicities, they are fully compatible within $1\sigma$, and (ii) at super-solar metallicity, the value of $f_{\rm CJ|SP}$ is higher than $f_{\rm CJ}$ by approximately $3/2$ ($\sim 30\%$ vs $\sim 20\%$; Fig. 1), but the significance of this difference is low, at $1.7\sigma$ (compared to $2.7\sigma$ in BL24 for $\rm[Fe/H]>0$).
Figure 1: Occurrence rates of CJs with $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ in the SP sample (red solid line) and the merged RV comparison sample (black dashed line) at sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicity. See Table 13.
Figure 2: Same as Fig. 1 for three different bins of stellar mass: $0.6-0.8\,\hbox{$\mathrm{M}_{\odot}$}$ (left), $0.8-1.0\,\hbox{$\mathrm{M}_{\odot}$}$ (center), and $1.0-1.2\,\hbox{$\mathrm{M}_{\odot}$}$ (right). See Table 14.
@@ 4.3 Frequency of CJs as a function of stellar metallicity and mass
We report $f_{\rm CJ|SP}$ and $f_{\rm CJ}$ for the same bins of stellar metallicity as above and for the three ranges of stellar mass discussed in Sect. 3, in Table 14. We show these values in Fig. 2 (see also Fig. 10 for $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$).
We find an excess of CJs in SP systems at $2.5\sigma$ significance level only for super-solar mass and metallicity, namely $M_{\star}\geq 1.0\,\hbox{$\mathrm{M}_{\odot}$}$ and $\rm[Fe/H]>0$ (see the right panel in Fig. 2, and Table 14), in agreement with the recent findings of BL25. This is the origin of the higher $f_{\rm CJ|SP}$ compared to $f_{\rm CJ}$ at super-solar metallicity and integrated stellar masses (Sect. 4.2 and Fig. 1).
@@ 4.4 SP-CJ relations as a function of SP and CJ multiplicity
We explored possible SP-CJ trends as a function of SP multiplicity. For this analysis, we adopted the $0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ mass range (instead of $0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$), given the higher number of systems with both CJs and SPs (27 compared to 23). If the presence of CJs hinders the formation or migration of SPs, one would expect a lower degree of SP multiplicity in the presence of CJs. For example, low-mass “jumpers” overtaking the CJ towards inner orbits are more likely to form single SP systems (see Fig. 3 in Izidoro et al. 2015). By comparing the frequency of multiple systems with a number of SPs $n_{\rm p}\geq 2$ in the SP subsamples with and without CJs, we find $f_{n_{\rm p}\geq 2,\rm w/\,CJ}=37^{+10}_{-8}\%$ (10 out of 27 systems) and $f_{n_{\rm p}\geq 2,\rm w/o\,CJ}=43^{+4}_{-3}\%$ (81 out of 190 systems), respectively. These rates are statistically consistent within uncertainties.
We also searched for possible (anti)correlations as a function of CJ multiplicity to investigate whether two or more CJs may hinder SP formation more than a single CJ system (Sect. 1). To this end, we compared the frequency of single SPs in systems with one CJ to that in systems with two (or more) CJs. The latter should be higher if two or more CJs have a more deleterious effect than a single CJ on the formation of multiple SP systems. We find no evidence for this: the rates of single SP systems with one CJ and two CJs are compatible, namely $63^{+9}_{-12}\%$ (12 out of 19 systems) and $62^{+13}_{-18}\%$ (5 out of 8 systems), respectively (see Fig. 7).
@@ 4.5 SP-CJ relation as a function of SP composition
We investigated the possible presence of an architecture-composition link by determining the frequency of rocky SPs [^10] We define a rocky planet as one located between the 100% iron and 100% silicate compositions in the mass-radius diagram of small planets (e.g., Zeng & Sasselov 2013). in CJ systems. This analysis tests whether the formation of rocky planets is favored in the presence of CJs if CJs block the migration of icy sub-Neptunes from beyond the water snowline (Schlecker et al. 2021). We derive $f_{\rm rocky,w/\,CJ}=33^{+13}_{-9}\%$ from five transiting systems with rocky planets (HD219134, Kepler-139, Kepler-407, K2-312, and K2-364) out of fifteen in total (including three RV systems that were later found transiting). This indicates that systems with both rocky planets and CJs do not represent the majority of the SP-CJ transiting systems.
@@ 4.6 SP multiplicity as a function of CJ eccentricity
When ordering the SP systems with CJs as a function of the CJ eccentricity (Fig. 3), only single or, more rarely, two SPs are found with eccentric ($e\gtrsim 0.4$) CJs, the most extreme case being K2-312/HD 80653 (B23). In contrast, higher multiplicity (up to $n_{\rm p}=4$ for HD 219134) SP systems are found in the company of CJs in quasi-circular orbits. This is consistent with the expectation of stronger dynamical perturbations at inner orbital separations by eccentric CJs (e.g., Huang et al. 2017).
Figure 3: Architectures of 27 systems with both SPs and CJs in the mass range $0.3\leq{\rm M_{p}}\leq 13\,\hbox{$\mathrm{M}_{\rm Jup}$}$, ordered according to increasing orbital eccentricity of the CJs (from top to bottom; for multiple CJs, the highest eccentricity is considered). The planet circles have sizes proportional to their mass (see legend in the top right corner), and colors indicating orbital eccentricity (more eccentric planets are shown in yellow). Host stars are color coded as a function of stellar metallicity (metallicity increases from blue to red). All systems with three or more inner SPs are accompanied by CJs with low eccentricities.
@@ 5 Discussion and conclusions
5.1 Cold Jupiters and small planets: friends, foes, or indifferent?
@@
In this work, we first investigated the possible SP-CJ relation as a function of stellar metallicity and mass. This did not reveal any correlation or anticorrelation (i.e., SPs and CJs are neither friends nor foes) in eight of the nine mass and metallicity bins, that is, at sub-solar mass for any metallicity and at super-solar mass for solar and sub-solar metallicity. We instead find an enhancement of CJs in SP systems (i.e., SPs and CJs could be friends) at super-solar mass and metallicity, though at less than $3\sigma$, in general agreement with BL25.
@@
We do not find any SP-CJ trends with planet multiplicity and composition, which further challenge the anticorrelation models predicting a dearth of CJs in SP systems (Sect. 1). Nonetheless, it should be noted that (i) the limited subsamples of SP systems with CJs (27 systems) inevitably result in large uncertainties in CJ occurrence rates, making it difficult to detect possible weak correlations or anticorrelations with planet properties; and (ii) the SP multiplicity considered in this work is based on the detected SPs and may be different from the true multiplicity, especially if non-transiting SPs with relatively long periods ($P\sim 25-100$ d) and/or low mass were not detected by RV follow-up.
@@
Our investigation of the architecture-composition relation in Sect. 4.5 must be considered with caution because it only concerned rocky planets. In principle, it may also apply to rocky planets with tenuous H/He atmospheres, if some available gas was accreted by the super-Earths at the time of their formation, and was not subsequently stripped away by post-formation processes such as atmospheric photoevaporation (e.g., Lopez & Fortney 2014) or core-powered mass loss (e.g., Ginzburg et al. 2018). However, the well-known degeneracy in SP composition from the measurement of the bulk density alone does not allow us to distinguish whether or which non-rocky planets in the ten transiting systems with CJs of our SP sample are actually dry (rocky worlds shrouded by thin H/He atmospheres, which formed within the water snowline) or icy (water worlds with possible H/He envelopes, formed beyond the water snowline and then migrated inwards). This makes the investigation of the architecture-composition link challenging.
@@
We show that higher multiplicity SP systems are found in systems with CJs in low-eccentricity orbits, being immune to dynamical instabilities generated by highly eccentric CJs. In this sense, SPs are not indifferent to CJs, even in the majority of cases where $f_{\rm CJ|SP}\sim f_{\rm CJ}$.
5.2 Caveats of this work, BL24, and BL25
@@
The occurrence rates derived in this work, BL24, and BL25 may be slightly underestimated due to two primary factors. However, these factors should affect both the SP and comparison samples approximately equally, making the search for correlations, in principle, reliable. Firstly, the completeness values computed in our work and in BL24 are optimistic because they only account for the formal RV uncertainties and not for uncorrelated jitter (as addressed comprehensively by B23), which in some cases may be considerably higher than the former [^11] Nevertheless, proper Markov chain Monte Carlo analyses of the RV datasets of $\sim 1350$ systems considered here, similarly to B23, would require unrealistic computation times.. Secondly, a small number of linear or quadratic RV trends, which might later result in CJ signals, were not taken into account in the computation of the occurrence rates, which is based only on confirmed CJs. Nevertheless, among the targets in the SP sample with relatively short temporal baselines, that is $1\leq\Delta T\leq 4$ yr, we identified only three targets with RV long-term trends still compatible with CJs (namely, TOI-1279, TOI-1742, and TOI-1778). Moreover, such trends are very often induced by brown dwarfs or low-mass stellar companions, or they correlate with activity indicators (see B23).
@@
Conversely, occurrence rates may decrease if the minimum masses ($M_{\rm p}\sin{i}$) of some of the considered CJs turn out to be true masses ($M_{\rm p}$) above the 20 (or 13) $\mathrm{M}_{\rm Jup}$ upper limit, corresponding to brown dwarfs or low-mass stellar objects, due to low orbital inclinations (e.g., Xiao et al. 2023).